London purchased a piece of real estate last year for $86,000. The real estate is now worth $104,600. If London needs to have a total return of 0.22 during the year, then what is the dollar amount of income that she needed to have to reach her objective? Round to two decimal places.
Answer #1

Total return==(End value-Beginning value+Income)/Beginning value  

0.22=(104600-86000+Income)/86000

(0.22*86000)=18600+Income

Income=18920-18600

=$320